OpenStudy (anonymous):

Hi. someone can explain to me how I can do this problems please. Tell whether the fraction is in simplest form. Write yes or no. You must show your work! 4/6 5/9 10/12 11/16

OpenStudy (anonymous):

\[no \frac{ 4 }{6 }\rightarrow \frac{ 2 }{ 3 } (divide by 2)\]

OpenStudy (anonymous):

yes for #2

OpenStudy (anonymous):

yes for last one

OpenStudy (anonymous):

\[no \frac{ 10 }{ 12 }\rightarrow \frac{ 5 }{ 6 }(divide by 2)\]

OpenStudy (anonymous):

Thans udtdtu and God bless you